One has the following information about the sector of companies: Direct tax: 1.6 times larger than the Distributed Profits, Distributed Profits: 15% of the Profits from production and patrimony (income), Current transfers toward other sectors: 3% from the Disposable Income, Disposable income: 61.3 billion Euro, Profits from production and patrimony (income): 79.5 billion Euro. Use an account of the sector of companies and compute the value of the current transfers received from other sectors!


1
Expert's answer
2022-01-23T15:40:37-0500

"61.3-1.6\times(79.5\times0.15)+(x-0.03*\times61.3)=79.5-79.5\times0.15"

x=27.194